OpenStudy (anonymous):

Because of friction and air resistance, each swing of a pendulum is a little shorter than the previous one. The lengths of the swings form a geometric sequence. Suppose the first swing of the pendulum has an arc length of 100 cm and a return swing of 99 cm. On which swing will the length first have a length less than 50 cm? Find the total distance traveled by the pendulum until it comes to rest.

OpenStudy (anonymous):

Im just stuck on this one problem.

jimthompson5910 (jim_thompson5910):

sounds like 100 is the first term of the geometric sequence and 99 is the second term

jimthompson5910 (jim_thompson5910):

if so, then common ratio = (second term)/(first term) r = (second term)/(first term) r = 99/100 r = 0.99

jimthompson5910 (jim_thompson5910):

Your nth term would then be an = a*r^(n-1) an = 100*0.99^(n-1)

jimthompson5910 (jim_thompson5910):

Now plug in an = 50 and solve for n to get your answer

OpenStudy (anonymous):

plug in 50? an= 100*0.99^50-1?

jimthompson5910 (jim_thompson5910):

no into an, not n

jimthompson5910 (jim_thompson5910):

an = 100*0.99^(n-1) 50 = 100*0.99^(n-1) now solve for n

OpenStudy (anonymous):

99^n-1?

OpenStudy (anonymous):

:( dang..

jimthompson5910 (jim_thompson5910):

50 = 100*0.99^(n-1) 50/100 = 0.99^(n-1) 0.5 = 0.99^(n-1) ln(0.5) = ln(0.99^(n-1)) ln(0.5) = (n-1)ln(0.99) ln(0.5) = n*ln(0.99)-ln(0.99) ln(0.5)+ln(0.99) = n*ln(0.99) ( ln(0.5)+ln(0.99) )/ln(0.99) = n n = ( ln(0.5)+ln(0.99) )/ln(0.99) n = 69.9675639365284 n = 70 (round up) So on the 70th swing (and up), the arc length will be less than 50